### v3

@@ Abstract
Understanding the spatiotemporal patterns of human mobility is crucial for addressing societal challenges, such as epidemic control and urban transportation optimization. Despite advancements in data collection, the complexity and scale of mobility data continue to pose significant analytical challenges. Existing methods often result in losing location-specific details and fail to fully capture the intricacies of human movement. This study proposes a two-step dimensionality reduction framework to overcome existing limitations. First, we construct a potential landscape of human flow from origin-destination (OD) matrices using combinatorial Hodge theory, preserving essential spatial and structural information while enabling an intuitive visualization of flow patterns. Second, we apply principal component analysis (PCA) to the potential landscape, systematically identifying major spatiotemporal patterns. By implementing this two-step reduction method, we reveal significant shifts during a pandemic, characterized by an overall declines in mobility and stark contrasts between weekdays and holidays. These findings underscore the effectiveness of our framework in uncovering complex mobility patterns and provide valuable insights into urban planning and public health interventions.
@@ 1 Introduction
Humans are inherently mobile beings who frequently travel from one location to another for various daily activities, such as commuting, shopping, entertainment, and education. These individual movements collectively shape the intricate dynamics of urban mobility. Spatiotemporal patterns of human flow are crucial for a wide range of applications, including controlling the spread of epidemics [1, 2, 3], understanding social and economic issues such as social segregation [4, 5, 6] and urban gender disparities [7, 8], and integrating public transportation systems with urban design plans to enhance mobility efficiency [9, 10, 11, 12].
@@
Human flow is not static but fluctuates dynamically over time. Therefore, understanding the evolution of these patterns is crucial. For example, on weekdays, cities exhibit distinct human flow patterns during morning and evening rush hours, whereas on weekends, movement trends differ significantly. Furthermore, external factors such as natural disasters and pandemics can drastically alter these patterns. Natural disasters, such as earthquakes or floods, can abruptly disrupt natural human movement, triggering mass evacuations or substantial changes in commuting behaviors. Similarly, the outbreak of a pandemic, such as Coronavirus Disease 2019 (COVID-19) [13], can cause unprecedented shifts in human mobility, with lockdowns and travel restrictions drastically reducing movement and reshaping human flow patterns on a global scale. These temporal variations in human flow result from a complex interplay of factors, including economic activity, social events, environmental conditions, and policy interventions. Examining these patterns over time provides valuable insights into the underlying mechanisms driving human mobility.
@@
Recent advancements in data collection and processing technologies have enabled the analysis of large-scale mobility data at high spatiotemporal resolutions, facilitating the exploration of dynamic mobility patterns and the identification of underlying structures in human flow. We live in an era where vast amounts of human mobility data are generated from various sources [14, 15, 16, 17, 18]. Traditionally, mobility data have been gathered through censuses and surveys that collect information on where people live and work. With the increasing prevalence of mobile phones, call detail records (CDRs) have become an important data source. These records, provided by mobile network operators, track the location of a signal tower when a phone call or text message is sent. Additionally, global positioning system (GPS) data collected from GPS-enabled devices and smartphones offer increasingly detailed location information. Social media platforms also contribute valuable mobility data through user interactions and geotagged posts.
@@
Despite the availability of vast amounts of data, understanding the spatiotemporal patterns of human flow remains challenging owning to the large volume and complexity of such data. In previous research, mobility data have often been reduced down to temporal trends of macroscopic statistics, such as displacement and waiting time distributions [19], radius of gyration [20], number of visited places and visitation frequency [21], predictability [22], motifs [23], and mobility synchronization [24]. Some studies have examined the dynamics of one or more of these metrics [25, 26, 27, 28, 29, 30, 31, 24], revealing common patterns in human mobility. However, these data processing approaches often mitigate information overload by reducing detail, resulting in the loss of location-specific information. In other instances, data have been transformed into spatial distributions of temporal populations [32, 33, 34], preserving location-specific details but losing crucial mobility information, such as trip distances and their origins. Additionally, data have been used to fit general models of human mobility, including the gravity model [35], the intervening opportunities model [36, 37, 38, 39, 40], the radiation model [41, 42, 43, 44, 45, 46, 47, 48, 49], and the exploration and preferential return (EPR) model [21, 50, 51, 52, 53, 54]. Although these models offer valuable insights, systematic examination of the dynamics of human mobility patterns across different contexts is limited. In summary, previous approaches have failed to retain the full complexity of human mobility data. Methods that aggregate data into macroscopic statistics often lose spatial specificity, whereas those emphasizing spatial distributions omit essential trip information, such as origins and distances. Furthermore, model-based approaches abstract real-world patterns to fit theoretical constructs, which may not fully capture the intricacies of human movement.
To address those existing issues, this study proposes a two-step reduction method for human mobility data, enabling the effective visualization of spatiotemporal patterns in human flow (Fig. 1). As shown in the middle panel of Fig. 1, the first step involves utilizing the concept of the potential landscape of human flow [55, 56, 57, 58, 59]. In our case, the potential landscape is generated through applying the combinatorial Hodge theory [60, 61, 62, 63] to the origin-destination (OD) matrix. This method offers several advantages such as significantly reducing data dimensionality by transforming an OD matrix with $N^{2}$ elements for $N$ regions of interest into a potential landscape with only $N$ elements. Despite the reduction, human flow can still be accurately reproduced [55]. Furthermore, the potential landscape quantifies the relational mobility information between locations as location-level statistics, which can be naturally visualized on a map. This quantification facilitates the comparison of mobility patterns across different spatial scales. Accordingly, this study investigates the temporal evolution of human flow patterns. In the second step, the spatiotemporal dynamics of the potential landscape are categorized into spatial and temporal components using principal component analysis (PCA). PCA is a widely used statistical technique for dimensionality reduction in data analysis. Previous research has employed PCA to analyze and interpret human mobility patterns [64, 65, 66, 67]. By applying PCA to the time series of the potential landscape, as shown in the right panel of Fig. 1, the temporal trajectory of the score on the principal component axes is extracted, representing the contribution of each spatial grid to the corresponding eigenvector $\mathbf{w}_{k}$. This study investigates the spatiotemporal patterns of human mobility by constructing a scalar potential field for human flow and analyzing its temporal evolution. The findings reveal significant shifts in mobility behavior during the pandemic, including decreased overall movement and high spatially-constrained mobility patterns, particularly during peak commute hours. Additionally, a clear difference is observed between weekday and holiday mobility, with holidays showing delayed peaks and less spatial dispersion, both before and during the pandemic.
@@
Figure 1: Methodology for examining spatiotemporal mobility patterns. (Left) Raw, hourly OD data derived from mobility datasets serve as the foundation of the analysis, capturing detailed flow patterns between spatial grids. (Middle) Hourly OD data transformed into a series of hourly potential landscapes. (Right) Principal component axes representing the contribution of each spatial grid to the corresponding eigenvector $\mathbf{w}_{k}$. Hourly dynamics are represented by the trajectory of score changes over time.
@@
The remainder of this paper is organized as follows: Section 2 describes the mobility dataset, which contains hourly OD information for both pre-COVID-19 and COVID-19 periods. Section 3 provides an overview of combinatorial Hodge theory (Sec. 3.1) and explains its application to human flow data (Sec. 3.2). In Sec. 4, we begin with the extraction of the scalar potential field of human flow through the unique decomposition of a given OD matrix using combinatorial Hodge theory. Then, PCA is applied to the potential landscape to investigate temporal variations in human flows within a 24-hour period across different scenarios, including weekday versus weekend trips and pre-COVID-19 versus during COVID-19 periods. Finally, Section 5 summarizes our key findings and concludes the paper. These results highlight the profound impact of COVID-19 on urban mobility and provide insights into human flow adaptations to external disruptions, thereby offering valuable implications for urban planning and public health policies.
@@ 2 Data
[…]
The “LocationMind xPop” data refer to human flow data collected from individual location information transmitted by smartphones with users’ consent, provided by NTT DOCOMO, INC. To ensure privacy, the data provider processed the data collectively and statistically, protecting any personally identifiable information. The original location data comprised GPS coordinates (latitude and longitude) recorded at a frequency as short as every 5 minutes and did not include any information that could directly identify individuals.
The background map layouts in Figs. 2 and 3 are based on data by © OpenStreetMap contributors, available under the Open Database License (ODbL). Map tiles by © CARTO. The results in Figs. 2, 3, 4, and 5 were obtained by using data “LocationMind xPop © LocationMind Inc” [68].
@@ 3.1 Combinatorial Hodge theory
To extract essential insights from human mobility data, we apply combinatorial Hodge theory [60, 61, 62], a mathematical framework that decomposes a given vector field into three orthogonal components: the gradient of a scalar potential field, the curl of a vector potential field, and a harmonic component. When used in the analysis of human mobility data, this approach provides a structured method for examining the OD matrix, which represents the movement of individuals between locations [55, 56]. By introducing a scalar potential field through this decomposition, we effectively visualize the spatial distribution of human flow, reveal latent mobility patterns, and gain a deeper understanding of the underlying dynamics.
The potential landscape is derived using combinatorial Hodge theory [60, 61, 62] on an undirected graph $G(V,E)$, where $V$ denotes a set of vertices and $E$ represents a set of edges. Edge flow $Y$ is assigned to each edge, with each element $Y_{ij}$ representing the flow from vertex $i$ to vertex $j$. The edge flow is skew-symmetric, i.e., $Y_{ij}=-Y_{ji}$. The combinatorial gradient, curl, and divergence are defined as follows [62]:
@@
where ker($\Delta_{1}$) = ker(curl) $\cap$ ker(div) and $\text{curl}^{*}$ is the adjoint operator of the curl.
Potential $s$ is defined as the solution to the weighted least-squares optimization problem given by:
@@
[…]
@@
where $\Delta_{0}$ is the graph Laplacian defined by:
@@
where $\dagger$ denotes the Moore-Penrose inverse.
@@ 3.2 Application of combinatorial Hodge theory to mobility flow
where the element $Y_{ij}$ denotes the net flow from $i$ to $j$, satisfying the necessary skew-symmetric condition in the combinatorial Hodge theory [60, 61, 62]. To define the potential using the optimization problem in Eq. 5, the regional difference in potential, $s_{j}-s_{i}$, is evaluated in comparison to the net flow $Y_{ij}$, but only for location pairs included in the edge set $E$. This evaluation is further weighted using $W_{ij}$, which reflects the significance of each pair in the analysis.
@@
The determination of edges $E$ and weights $W_{ij}$ is strongly influenced by the distances between locations. To explore the dependence of mobility flow on distance, we analyze the percentage of non-zero mobility pairs ($M_{ij}>0$ or $M_{ji}>0$) as a function of road distance, using data from the Japanese mobility dataset (Sec. 2). The results are presented in the Supplementary Information. Our analysis revealed a sharp decline in the percentage of non-zero mobility pairs as the distance increases. Notably, even at a relatively short distance of approximately $d_{ij}\sim 40$ km, nearly half of the mobility pairs exhibit zero flow volume.
@@
The phenomenon of zero-mobility flow offers several important implications. Mobility incurs substantial costs, and zero flow suggests that the costs associated with travel outweigh the driving forces behind regional differences. At shorter distances, where travel costs are generally low, zero mobility flow between two locations implies that the regional difference $(s_{j}-s_{i})$ is minimal. Conversely, at longer distances, where costs become disproportionately high, the absence of mobility flow conveys little useful information about regional potential differences.
To address this issue, we apply a weighting method to systematically exclude trips that occur at distances where mobility is rarely observed. This method, which has been proven to perform well in urban settings [56], enabling us to define a threshold distance, denoted as $\theta$. The threshold $\theta$ is established as the distance below which the majority of trips occur, based on the probabilistic criterion $p(X<\theta)=0.99$ (Supplementary Information). Specifically, the weight $W_{ij}$ is expressed as:
$W_{ij}=\begin{cases}1\quad&\text{mobility pair $\{(i,j)\mid d_{ij}>\theta\}$}\\ 0\quad&\text{mobility pair $\{(i,j)\mid d_{ij}<\theta\}$}\end{cases}.$ (10)
For the set of edges $E$, we select pairs with positive weights, $\{(i,j)\mid W_{ij}=1\}$.
By applying combinatorial Hodge theory to the OD matrix, we transform the data as illustrated in Fig. 1, progressing from the left to the middle panel. Specifically, the OD information at each time step is decomposed into a one-dimensional scalar field, referred to as a potential landscape [55, 56]. This potential landscape serves as a compact yet powerful representation of spatial dynamics, preserving essential spatial information while reducing data dimensionality. Each spatial grid is assigned a unique potential value that encapsulates the relative attractiveness or activity level of the location. Notably, the potential landscape retains the key trip information encoded in the OD matrix. This preservation is achieved because the difference between potential values at two locations determines the likelihood and direction of trips. Individuals tend to move from areas of lower potential to areas of higher potential, aligning with the intuitive interpretation of flows driven by regional differences in potential.

### v4

@@ Abstract
Understanding human mobility is vital to solving societal challenges, such as epidemic control and urban transportation optimization. Recent advancements in data collection now enable the exploration of dynamic mobility patterns in human flow. However, the vast volume and complexity of mobility data make it difficult to interpret spatiotemporal patterns directly, necessitating effective information reduction. The core challenge is to balance data simplification with information preservation: methods must retain location-specific information about human flows from origins to destinations while reducing the data to a comprehensible level. This study proposes a two-step dimensionality reduction framework: First, combinatorial Hodge theory is applied to the given origin–destination (OD) matrices with timestamps to construct a set of potential landscapes of human flow, preserving imbalanced trip information between locations. Second, principal component analysis (PCA) expresses the time series of potential landscapes as a linear combination of a few static spatial components, with their coefficients representing temporal variations. The framework systematically decouples the spatial and temporal components of the given data. By implementing this two-step reduction method, we reveal large weight variations during a pandemic, characterized by an overall decline in mobility and stark contrasts between weekdays and holidays. These findings demonstrate the effectiveness of our framework in uncovering complex mobility patterns and its potential to inform urban planning and public health interventions.
@@ 1 Introduction
Humans are inherently mobile beings who frequently travel from one location to another for various daily activities, such as commuting, shopping, entertainment, and education. These individual movements collectively shape the intricate dynamics of urban mobility. Determining the spatiotemporal patterns of human flow are crucial for a wide range of applications, including controlling epidemics spread [JLY2020, VBS2006, HSA2021], understanding social and economic issues such as social segregation [AKK2014, MCD2021, YPM2023] and urban gender disparities [GTP2020, CDG2023], and integrating public transportation systems with urban design to enhance mobility efficiency [VSR2018, XCG2021, BZK2021, GPB2016].
@@
Human flow is not static but fluctuates dynamically over time. Therefore, understanding the evolution of these patterns is crucial. For example, on weekdays, cities exhibit distinct human flow patterns during morning and evening rush hours, whereas on weekends, movement trends differ significantly. Furthermore, external factors such as natural disasters and pandemics can drastically alter these patterns. Natural disasters, such as earthquakes or floods, can abruptly disrupt natural human movement, triggering mass evacuations or substantial changes in commuting behaviors. Similarly, the outbreak of a pandemic, such as Coronavirus Disease 2019 (COVID-19) [ZZW2020], can cause unprecedented shifts in human mobility, with lockdowns and travel restrictions drastically reducing movement and reshaping human flow patterns on a global scale. These temporal variations in human flow result from a complex interplay of factors, including economic activity, social events, environmental conditions, and policy interventions. Examining these patterns over time provides valuable insights into the underlying mechanisms driving human mobility.
@@
Recent advancements in data collection and processing technologies have enabled the analysis of large-scale mobility data at high spatiotemporal resolutions, facilitating the exploration of dynamic mobility patterns and the identification of underlying structures in human flow. We live in an era where vast amounts of human mobility data are generated from various sources [BDK2015, Z2015, ZTL2016, ZYW2018, YJR2022]. Traditionally, mobility data have been gathered through censuses and surveys that collect information on where people live and work. With the increasing prevalence of mobile phones, call detail records (CDRs) have become an important data source. These records, provided by mobile network operators, track the location of a signal tower when a phone call or text message is sent. Additionally, global positioning system (GPS) data collected from GPS-enabled devices and smartphones offers increasingly detailed location information. Social media platforms also contribute valuable mobility data through user interactions and geotagged posts.
@@
Although vast amounts of mobility data are now available, their sheer volume and complexity make it difficult to understand the spatiotemporal patterns of human flow on the Earth’s surface, necessitating effective information reduction methods to avoid information overload. Table 1 summarizes the information reduction methods in previous studies. In earlier research, mobility data have often been reduced to temporal trends of macroscopic statistics, such as displacement and waiting time distributions [BHG2006], radius of gyration [GHB2008], number of visited places and visitation frequency [SKW2010], predictability [SQB2010], motifs [SBC2013], and mobility synchronization [SBB2023]. Some studies have examined the dynamics of one or more of these metrics [KSZ2020, HSU2013, LBH2012, GY2017, WT2016, HRH2020, KSH2022, SBB2023], revealing common patterns in human mobility. Nevertheless, these data processing approaches often mitigate information overload by reducing detail, resulting in the loss of location-specific information, which prevents discussion of mobility patterns at specific locations on a map. In other instances, data have been transformed into spatial distributions of temporal populations [BLT2011, LXM2016, LST2017], preserving location-specific details but losing crucial mobility information, such as trip distances and their origins. Additionally, data have been used to fit general models of human mobility, including the gravity model [Z1946], intervening opportunities model [S1940, FM1990, AP2001, NSL2012, SYB2015], radiation model [SGM2012, LZD2013, YZF2014, REW2014, KLG2015, VTN2018, LY2019, LY2020, DF2022], and exploration and preferential return (EPR) model [PSR2015, SDO2021, BLE2015, ASS2018, LZZ2021]. Although these models offer valuable insights, the fitted parameters are usually macroscopic, and location-specific information is often lost. In summary, the core challenge is to reconcile data simplification with information preservation: To study mobility patterns on a map, retaining location-specific information about human flows from origins to destinations is crucial. At the same time, the data must be simplified sufficiently to be comprehensible.
Table 1: Comparison of information reduction methods of human mobility data
Method Captured properties Advantages Limitations
Fit information to general models of human mobility (e.g., gravity, radiation, intervening opportunities, EPR models) [Z1946, S1940, FM1990, AP2001, NSL2012, SYB2015, SGM2012, LZD2013, YZF2014, REW2014, KLG2015, VTN2018, LY2019, LY2020, DF2022, PSR2015, SDO2021, BLE2015, ASS2018, LZZ2021] Calibrated model parameters to given datasets Abstract real-world patterns to fit theoretical constructs. Often loss of location-specific information if the model has no such information
Reduce information to temporal trends of macroscopic statistics [BHG2006, GHB2008, SKW2010, SQB2010, SBC2013, SBB2023, KSZ2020, HSU2013, LBH2012, GY2017, WT2016, HRH2020, KSH2022] Temporal trends of macroscopic statistics (e.g., displacement, radius of gyration, predictability, motifs, synchronization) Effectively reduces the complexity and scale of mobility data, and is helpful in revealing common patterns in humans mobility and examining the dynamics of metrics over time. Loss of location-specific information
Transform information into spatial distributions of temporal populations [BLT2011, LXM2016, LST2017] Time series of population distributions Preserves location-specific details of population distribution Loss of crucial mobility information, such as trip distances and their origins
Two-step dimensionality reduction (Our method) Potential landscape as static spatial components and a trajectory of time-varying coefficients of these components Preserves essential location-specific spatial patterns of human mobility and their temporal changes Loss of circular components of human flow, which is orthogonal to the gradient component derived from potential
Therefore, our goal is to balance simplicity with information fidelity to study human flow on a map using information reduction methods. This study proposes a two-step reduction method for human mobility data, enabling the effective representation of spatiotemporal patterns in human flow. Fig. 1 illustrates the concept of our method. Let us consider a human flow data as a time series of origin–destination (OD) matrices (left panel in Fig. 1). The first reduction step involves utilizing the concept of the scalar potential [AFF2022, AFF2023, MMB2019, SOT2022, YLG2024]. In our case, the potential is generated through applying the combinatorial Hodge theory [K1949, H1989, JLY2011, BBS2025] to the OD matrix at each timestamp. This method has key advantages. (1) It retains location-specific information about human flows from origins to destinations. Higher-potential locations collect the people’s movements from lower-potential locations. It extracts the imbalanced component of a given flow, which can be expressed by the differences of potentials between locations (see Fig. 2 and Methods for details). (2) It offers a significant reduction in data dimensionality. An OD matrix represents the trips from origins to destinations. It signifies relational data between locations (i.e., locations $\times$ locations dimensionality), and is not easily shown on a map. By reducing the OD matrix to the potential at each location, the obtained potential landscape is a location-level statistic, which can be naturally visualized on a map. After the first step, we get a time series of potential landscapes (middle panel in Fig. 1), like time slices of a movie of a map. The second step decouples the temporal and spatial components. Using principal component analysis (PCA), the time series of potential landscapes can be approximated by linear combinations of a few potential landscapes. The potential landscapes (eigenvector $\mathbf{w}_{k}$) are static and only retain spatial information of human mobility. The mixing of these potential landscapes changes with time and the coefficients (principal component score) retain the temporal information of human mobility. As a result of the two-step reduction, spatiotemporal behavior in human flow can be represented as a trajectory in a low-dimensional space, as shown in the right panel of Fig. 1. Compared with the raw OD matrices, this two-step reduction generates a small number of potential landscapes (corresponding to the PCA axes) that provide reduced, location-specific flow information. In contrast, the resulting trajectory offers a clear representation of temporal changes.
As a case study, we apply the reduction method to study a dataset in Tokyo before and during COVID-19. Do the first dominant principal components capture weekday/holiday and pandemic temporal signatures? The findings reveal significant shifts in mobility behavior during the pandemic, including decreased overall movement and high spatially constrained mobility patterns, particularly during peak commute hours. Additionally, a clear difference is observed between weekday and holiday mobility, with holidays showing delayed peaks and less spatial dispersion, both before and during the pandemic.
@@
Figure 1: Methodology for examining spatiotemporal mobility patterns. (Left) Raw, hourly OD data derived from mobility datasets serve as the foundation of the analysis, capturing detailed flow patterns between spatial grids. (Middle) Hourly OD data transformed into a series of hourly potential landscapes. (Right) Principal component axes representing the contribution of each spatial grid to the corresponding eigenvector $\mathbf{w}_{k}$. Hourly dynamics are defined by the trajectory of score changes over time.
@@
The remainder of this paper is organized as follows: Section 2 describes the mobility dataset, which contains hourly OD information for both pre-COVID-19 and COVID-19 periods. Section 3 provides an overview of combinatorial Hodge theory (Section 3.1) and explains its application to human flow data (Section 3.2). In Section 4, we begin with the extraction of the scalar potential field of human flow through the unique decomposition of a given OD matrix using combinatorial Hodge theory. Then, PCA is applied to the potential landscape to investigate temporal variations in human flows within 24 hours across different scenarios, including weekday versus weekend trips and pre-COVID-19 versus during COVID-19 periods. Finally, Section 5 summarizes our key findings and concludes the paper. These results highlight the profound impact of COVID-19 on urban mobility and provide insights into human flow adaptations to external disruptions, thereby offering valuable implications for urban planning and public health policies.
@@ 2 Data
For this study, we use Japanese human mobility data, “LocationMind xPop,” provided by LocationMind Inc. [LocationMind], that encompasses GPS records collected within the Tokyo metropolitan area. The dataset spans two time periods: May and June 2019 (pre-COVID-19) and May and June 2021 (during COVID-19). These data include estimated home locations, origins, and destinations, which are mapped using a 2km by 2km spatial grid system based on the Japanese mesh index [MESH]. The dataset distinguishes between weekday and holiday movements and records the specific hours during which these movements occur. Notably, the data are aggregated, reflecting average flows over multiple days rather than individual trips. Additionally, the number of trips is based on the unique users recorded, with adjustments against population census totals, providing an estimate of real population movements.
The “LocationMind xPop” data refer to human flow data derived from individual location records transmitted by smartphones with users’ consent, provided by NTT DOCOMO, INC. To ensure privacy, the data provider processed the data in aggregate using statistical methods, protecting any personally identifiable information. The original location data comprised GPS coordinates (latitude and longitude) recorded at intervals as short as 5 minutes and did not include any information that could directly identify individuals.
We outline several limitations of the dataset. First, we relied on aggregated two-month mobility averages for 2019 and 2021, which limited detection of transient fluctuations, such as those driven by weather, short-term events, or sudden policy shifts. Second, we adjusted the GPS data using population census data, assuming proportional relationships between the sampled device users and the broader population. This assumption may overlook variations in local population density and disparities in device usage, thereby introducing bias.
The background map layouts in Figs. 4 and 5 are based on data from © OpenStreetMap contributors, available under the Open Database License (ODbL). Map tiles are provided by © CARTO. The results in Figs. 4, 5, 7, and 8 were obtained by using data from “LocationMind xPop © LocationMind Inc” [LocationMind].
@@ 3.1 Combinatorial Hodge theory
Figure 2: Illustration of potential extraction by combinatorial Hodge decomposition. Input edge flow $Y$ in the left panel can be decomposed into three orthogonal components: gradient, harmonic, and curl flows. The gradient flow is the focal component of this reduction method, which can be described by differences in node potentials, denoted by red numbers. The other elements are cyclic, and incoming and outgoing fluxes are balanced, while the gradient component represents the imbalanced flow that induces the temporal changes of population mass.
As the first step of reduction, we apply combinatorial Hodge theory [JLY2011], a mathematical framework that extracts the gradient component of a given flow from the other cyclic components. The gradient component represents the imbalanced flow of the given flow that induces the temporal changes of population mass: it can be derived from the differences of “potentials” at each location, and we refer to the geospatial distribution of the potentials as potential landscape [AFF2022, AFF2023]. We favor Hodge gradients because they summarize net source–sink structure interpretable on maps. This potential landscape provides a dimensionality reduction of human flow. The human flow—the trips from origins to destinations—represents relational data between locations (i.e., locations $\times$ locations dimensionality), and the reduced potential landscape is a location-level statistic.
To describe the potential landscape mathematically, we consider an undirected graph $G(V,E)$, where $V$ denotes a set of vertices (corresponding to locations) and $E$ represents a set of edges. Edge flow $Y$ is assigned to each edge, with each element $Y_{ij}$ representing the flow from vertex $i$ to vertex $j$. The edge flow is skew-symmetric, i.e., $Y_{ij}=-Y_{ji}$. The combinatorial gradient, curl, and divergence are defined as follows [JLY2011]:
@@
where ker($\Delta_{1}$) = ker(curl) $\cap$ ker(div) and $\text{curl}^{*}$ is the adjoint operator of curl.
Potential $s$ is obtained by solving the weighted least-squares optimization problem given by
@@
where $W_{ij}(\in[0,1])$ is the weight for each pairwise comparison, primarily determined by the distance $d_{ij}$ as discussed in Section 3.2. The optimization problem determines the closest point to the given data $Y$ in the subspace of edge flows, which comprises the gradient flows of the potential. The problem can be solved by an $l_{2}$-projection of $Y$ onto $\text{im}(\text{grad})$. With a Euclidean inner product in space $\mathcal{Y}$, $\langle X,Y\rangle=\sum_{\{i,j\}\in E}W_{ij}X_{ij}Y_{ij}$, the normal equation can be expressed as:
@@
where $\Delta_{0}$ is the graph Laplacian given by
@@
where $\dagger$ denotes the Moore–Penrose inverse.
@@ 3.2 Application of combinatorial Hodge theory to mobility flow
where the element $Y_{ij}$ indicates the net flow from $i$ to $j$, satisfying the necessary skew-symmetric condition in the combinatorial Hodge theory [K1949, H1989, JLY2011]. To define the potential using the optimization problem in Eq. 5, the regional difference in potential, $s_{j}-s_{i}$, is compared with the net flow $Y_{ij}$, but only for location pairs included in the edge set $E$. This evaluation is further weighted using $W_{ij}$, which encodes the importance of each pair in the analysis.
@@
The determination of edges $E$ and weights $W_{ij}$ depends primary on the distances between locations. To explore the dependence of mobility flow on distance, we analyze the share of nonzero mobility pairs ($M_{ij}>0$ or $M_{ji}>0$) as a function of road distance, using data from the Japanese mobility dataset (Section 2). The results appear in the Supplementary Information. Our analysis reveals a sharp decline in the percentage of non-zero mobility pairs as the distance increases. Notably, even at a relatively short distance of approximately $d_{ij}\sim 40$ km, nearly half of the mobility pairs exhibit zero flow volume.
@@
The phenomenon of zero-mobility flow admits several interpretations. A possible explanation is that mobility incurs substantial costs, and zero flow suggests that the costs associated with travel outweigh the drivers of regional differences. If the high costs prevent many people from traveling, the absence of mobility flow conveys little useful information about regional potential differences, and it will be better to ignore it in the optimization by (5). At shorter distances, where travel costs will be generally low, zero mobility flow between two locations implies that the regional difference $(s_{j}-s_{i})$ is too minimal to induce travel. In this situation, the absence of mobility flow provides valuable information and should be considered in the optimization in (5). Importantly, while we cannot identify the specific reasons preventing travel, if trips are consistently unobserved, this zero-flow pair conveys no information about regional potential differences and is therefore excluded from the analysis.
To exclude the zero mobility flow with little information, we apply a weighting method with a threshold distance $\theta$, which has been shown to perform well in urban settings [AFF2023]:
$W_{ij}=\begin{cases}1\quad&\text{mobility pair $\{(i,j)\mid d_{ij}\leq\theta\}$}\\ 0\quad&\text{mobility pair $\{(i,j)\mid d_{ij}>\theta\}$}\end{cases}.$ (10)
The threshold $\theta$ is the distance below which the majority of trips occur, using the probabilistic criterion $p(X<\theta)=0.99$ (Fig. 3). The value ($p$ = 0.99) is chosen to include the majority of trips while excluding possible outliers with less than one percent. We confirmed that the obtained results by this specific criterion with $p=0.99$ are robust against small perturbations on $p$ (see Supplementary Information for details). For the set of edges $E$, we retain pairs with positive weights, $\{(i,j)\mid W_{ij}=1\}$.
Figure 3: Threshold distances $\theta$ across four scenarios: (a) 2019 Weekday; (b) 2019 Holiday; (c) 2021 Weekday; and (d) 2021 Holiday.